OpenStudy (anonymous):

The half-life of Radium-226 is 1590 years. If a sample contains 100 mg, how many mg will remain after 2000 years?

OpenStudy (anonymous):

half-life formula: m(t) = m(sub o) e^(-rt) where r = ln2/h m(t) = 100 * e^[-(ln2/1590)*2000] m(t) = 100 * e^(-.87188) m(t) = 100 * .418165 m(t) = 41.8165 mg

OpenStudy (anonymous):

appreciate the help!